2 C2H6(g)      +      2 O2(g) =  CO2(g)    +      2 H2O(I)

1. How many grams of H2O will be produced by 58.2 L of CH4 at STP? Assume an excess of O2.


1
Expert's answer
2021-10-18T10:49:39-0400

We use the fact that 1 mole of any gas at STP has a volume of 22.4 L to find how much water was produced:


"\text{58.2 L of CH}_4 \text{ at STP} \times\cfrac{\text{1 mol of CH}_4}{\text{22.4 L of CH}_4}\times\cfrac{\text{2 mol of H}_2O}{\text{1 mol of CH}_4}\times\cfrac{\text{18 g of H}_2O}{\text{1 mol of H}_2O}
\\ =93.54\text{ g of H}_2O"


In conclusion, 93.54 g of water was produced during the combustion.



Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.